OpenStudy (anonymous):

Find the coordinates of the vertices and foci of the hyperbola ((y-1)^2)) /25 - (x+2)^2 = 1 HELP :(

OpenStudy (anonymous):

\[\frac{ (y-1)^2 }{ 25 } - (x+2)^2 = 1\]

OpenStudy (anonymous):

im only in 6th grade XD no lie so im going to do the math, brb

OpenStudy (mathmale):

I'd suggest you re-write the original equation as follows:

OpenStudy (anonymous):

hmmmm, its not { y-1} 2^

OpenStudy (anonymous):

is it?

OpenStudy (anonymous):

what do you mean Jazzi?

OpenStudy (mathmale):

\[\frac{ (y-1)^2 }{5^{2} } - \frac{ (x+2)^2 }{ 1^{2} }= 1\]

OpenStudy (anonymous):

To complete the problem I need to find dy/dx at (1, -1) so I substitute x = 1, y = -1 into the equation above to get... wait u said cooranents hmm brb again lol

OpenStudy (mathmale):

Compare this equation to the one you typed in originally.

OpenStudy (mathmale):

@JazziLeeRiser : This problem does not involve calculus.

OpenStudy (anonymous):

@mathmale you got it,

OpenStudy (anonymous):

I get what you did now :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

im going to medel u @mathmale

OpenStudy (mathmale):

OK, good. We can get a lot of info from this re-written equation for your hyperbola. Can you look at it and determine the CENTER of y our hyperbola? Can you use the 5 and the 1 from this re-written equation to draw a rectangular box around the center of your hyperbola?

OpenStudy (anonymous):

medial*

OpenStudy (anonymous):

Sorry, I don't think I can get the center from the equation. Can you please help me?

OpenStudy (mathmale):

Your equation involves (y-1)^2. Thus, the y-coordinate of your center is 1.

OpenStudy (anonymous):

is it 1, -2 ?

OpenStudy (mathmale):

Your equation involves (x+2)^2. Thus, the x-coord of the center is ... what?

OpenStudy (mathmale):

Remember that when you write a point as (a,b), the x-coord comes FIRST and the y-coord comes SECOND. Try again. The center of your hyperbola is ... ?

OpenStudy (anonymous):

That's right, sorry! (-2,1)

OpenStudy (mathmale):

Right you are. Now, what information does that 5^2 provide you?

OpenStudy (anonymous):

it gives us a^2 ad we need that with the equation c^2=a+b^2 right?

OpenStudy (mathmale):

Let me change your wording a bit: "It gives us the value of a, which we will insert into the equation c = a + b." Better check this out on the Internet or in your textbook! There are NO squares here.

OpenStudy (mathmale):

a = 5. Yes, I agree. What is b? b = ? Please refer to the following: http://www.purplemath.com/modules/hyperbola.htm

OpenStudy (anonymous):

I was looking at the foci equation (c^2=a^2+b^2) my bad!

OpenStudy (mathmale):

I was wrong and you were mostly correct: You do need to use a^2, b^2 and c^2. Again, please refer to http://www.purplemath.com/modules/hyperbola.htm

OpenStudy (anonymous):

so doesn't a= + - square root of 5 ?

OpenStudy (mathmale):

Notice that in http://www.purplemath.com/modules/hyperbola.htm a, b and c are illustrated. Also note that a, b and c as showin in that diagram are positive. You are thinking of 25, I believe; the sqrt of 25 is 5. But we've already been talking in terms of 5. No need to find the square root of 5.

OpenStudy (anonymous):

oh right!

OpenStudy (anonymous):

b=1 right?

OpenStudy (mathmale):

Referring to the diagram in http://www.purplemath.com/modules/hyperbola.htm , a is measured horizontally and is the distance between the center and the vertex. Can you determine the value of a? b is measured vertically and is the distance between the center and the vertex. Can you determine b? Once you have a and b, you can find c.

OpenStudy (mathmale):

The diagram shows that a is measured horizontally in this particular problem; therefore, a=1. b is measured vertically and here is 5. Agree or disagree?

OpenStudy (mathmale):

Can you now draw y our hyperbola following the example in http://www.purplemath.com/modules/hyperbola.htm ? Show a=1 and b=5.

OpenStudy (anonymous):

I am confused on why a=1 if 25 is a ^2 ?

OpenStudy (mathmale):

Use the Draw utility.

OpenStudy (anonymous):

I don't need to graph it, I just need to find the coordinates and foci :(

OpenStudy (mathmale):

Again, please refer to the example in http://www.purplemath.com/modules/hyperbola.htm. Is a measured vertically or horiz? Is b measured vertically or horiz?

OpenStudy (mathmale):

True, you don't need to graph it, but doing so will make it very obvious where your vertices are.

OpenStudy (anonymous):

Oh ok, sorry!

OpenStudy (mathmale):

Again, the center is at (-2, 1). If a = 1, and a is a horizontal distance, then what are the coordinates of the right vertex? Next, find the coordinates of the left vertex.

OpenStudy (mathmale):

I'm stopping here because OpenStudy says you are offline.